OpenStudy (doc.brown):

I'm trying to clean this up, get (x-1) to be exposed.

OpenStudy (doc.brown):

\[(sqrt(x-1)-(sqrt(2))^2\]

OpenStudy (doc.brown):

\[(\sqrt(x-1)-\sqrt(2))^2\]

OpenStudy (sdfgsdfgs):

u can use this: (a-b)^2 = a^2 - 2ab + b^2